My question is off a worksheet that if called Solving problems involving inequalities. It is "Find the largest pair of consectutive integers whos sum is less than 85."
My work is:
x+x+2<85
2x<83

Found 2 solutions by checkley75, stanbon:
Answer by checkley75(3666)

X+X+2<85
2X<85-2
2X<83
X<83/2
X<41.5
PROOF SET X=41 & X+2=41+2=43
41+43<86
85<86 STATEMENT IS TRUE.

Answer by stanbon(75887)

"Find the largest pair of consectutive integers whose sum is less than 85."
------
Let the consecutive integers be "x" and "x+1"
EQUATION:
x+x+1<85
2x < 84
x < 42
x=41 and x+1=42
============
Cheers,
Stan H.